A two-digit number is six times the sum of its digits. The tens digit is 1 more than the units digit. Find the number. I dont even know where to begin with this one. Thanks for the help
Found 2 solutions by Paul, checkley71:
Answer by Paul(988)

Tens digit = x
One's digit = y
x=1+y (subsitution)
10x+y=6(x+y)
10x+y=6x+6y
Subsitute for x:
10(1+y)+y=6(1+y)+6y
10+10y+y=6+6y+6y
10+11y=6+12y
y=4

x=1+4
x=5

Hence, the number is 54 with 5 as tens digit and 4 as ones digit.
Paul.

Answer by checkley71(8403)

10X+Y=6(X+Y) & X=Y+1 SUBSTITUTION Y+1 FOR X IN THE FIRST EQUATION WE GET
10(Y+1)+Y=6((Y+1)+Y) OR 10Y+10+Y=6Y+6+6Y OR 11Y+10=12Y+6 OR -Y=-4 OR Y=4
THEREFORE X=4+1 OR X=5 THUS THE NUMBER IS 54 BECAUSE 5+4=9 AND 54=6 TIMES 9